OpenStudy (anonymous):

Write an equation in slope-intercept form that passes though thee point (3,-9)

OpenStudy (anonymous):

is there a second point?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It has the slope of -5

OpenStudy (anonymous):

well then, for fun you can set the slope to be "-3" y=mx+b we will need to find a proper b then given that the point passes though (3, -9) and that we will set the slope m as -3 now solve for b that will give you y= -3x + ____ and this will be one possible answer

OpenStudy (anonymous):

oh, you know the slope... that is slightly different you do the same thing, but plug in your proper slope rather than my -3 and you will again solve for b.. This will give you a slightly more interesting b than my example above but the mechanics are the same.

OpenStudy (anonymous):

if you know the slope is 5 plug it into \(y=mx+b\) along with the given x and y and solve for b once you find b plug it and the slope into the formula and that is its slope intercept form

OpenStudy (anonymous):

Wait How Do I Find B This Is All New To me

OpenStudy (anonymous):

plug in the values:\[-9=-5(3)+b\]multiply:\[-9=-15+b\]add 15 to both sides:\[4=b\]plug in the slope and b to the formula:\[\large{y=-5x+4}\]

OpenStudy (anonymous):

and that would be your answer @TajaAmari